### v2

# Normalizing Flows are Capable Models for RL

Abstract
Modern reinforcement learning (RL) algorithms have found success by using powerful probabilistic models, such as transformers, energy-based models, and diffusion/flow-based models. To this end, RL researchers often choose to pay the price of accommodating these models into their algorithms – diffusion models are expressive, but are computationally intensive due to their reliance on solving differential equations, while autoregressive transformer models are scalable but typically require learning discrete representations. Normalizing flows (NFs), by contrast, seem to provide an appealing alternative, as they enable likelihoods and sampling without solving differential equations or autoregressive architectures. However, their potential in RL has received limited attention, partly due to the prevailing belief that normalizing flows lack sufficient expressivity. We show that this is not the case. Building on recent work in NFs, we propose a single NF architecture which integrates seamlessly into RL algorithms, serving as a policy, Q-function, and occupancy measure. Our approach leads to much simpler algorithms, and achieves higher performance in imitation learning, offline, goal conditioned RL and unsupervised RL. [^1] Project page and code - https://rajghugare19.github.io/nf4rl/.
Figure 1: Probabilistic models and RL. Three capabilities of probabilistic models important for RL: efficient likelihood computation and sampling, and compatibility with variational inference (VI). These three capabilities are at the core of most RL algorithms and enable myriad applications in RL (See Section 4 for details). For example, behavior cloning requires maximizing likelihoods and sampling [4], GCRL value functions require estimating future discounted state probabilities [22], and MaxEnt RL and RL finetuning requires variational inference [53]. However, models like diffusion / flow matching and autoregressive transformers used in RL today support only a subset of these capabilities, making their application complicated. In this paper, we show how normalizing flow models, which posses these properties, can be a generally capable model family for RL
Table 1: Capabilities of normalizing flows and other probabilistic models. NFs have capabilities that make them applicable to a large number of algorithms in many RL problems (Section 4), including the ability to compute exact likelihoods, to sample from complex distributions, and to be trained via both maximum likelihood and variational inference (Section 3). Despite these excellent capabilities, research on NFs in RL has been limited compared to other state of the art models. Is this because NFs lack expressivity? Our experiments (Section 6) provide compelling evidence that this is not the case.
Model family Exact Likelihoods Sampling Variational Inference*
Variational Autoencoders Lower bound Yes No
Generative Adversarial networks No Yes No
Energy Based models Unnormalized MCMC No
Diffusion models Yes, via ODEs Unrolling SDEs No
Normalizing Flows Yes Yes Yes
    Although directly performing VI (posterior sampling and estimation) is generally intractible for the first four model families we have still marked them with yellow because additional algorithms for VI do exist [8].
1 Introduction
Reinforcement learning (RL) algorithms often require probabilistic modeling of high dimensional and complex distributions. For example, offline RL or behavior cloning (BC) from large scale robotics datasets requires learning a multi-modal policy [4, 15, 7] and value function estimation in goal-conditioned RL (GCRL) requires estimating probabilities over future goals [22]. Different algorithms require different capabilities, including likelihood estimation, variational inference (VI) and efficient sampling (see Figure 1). A model trained via BC must maximize action likelihoods during training and support fast action sampling for real-time control. Most offline or finetuning RL algorithms require policy models to be able to backpropagate gradients of the value function through sampled actions. For MaxEnt RL with continuous actions, models should be optimizable using variational inference. However, obtaining these abilities in current models often comes at the cost of compute and complexity (Table 1). For example, diffusion models [76] are highly expressive yet computationally demanding: training, sampling, and evaluating log likelihoods all require solving a differential equation. Autoregressive transformer models [84] are scalable and expressive, but are typically limited to discrete settings; using them for RL in continuous spaces requires learning discrete representations of actions [74] or states [59]. Gaussian models offer computational efficiency but lack expressivity. Can we design expressive probabilistic models which all the capabilities in Fig. 1?
Normalizing Flows (NFs) are among the most flexible probabilistic models: they $(1)$ support both likelihood maximization and variational inference training, and $(2)$ enable efficient sampling and exact likelihood computation (see Table 1). Yet, NFs have received far less attention from the RL community, perhaps due to the (mis)conception that they have restricted architectures or training instabilities [64]. This paper demonstrates that neither of these notions hold NFs back in practice: NFs can match the expressivity of diffusion and autoregressive models while being significantly simpler. Moreover, unlike other models, NFs do not require additional machinery to be integrated with existing RL frameworks, enabling simpler algorithms that are competitive with if not better than their counterparts.
Our primary contribution is to shed light on the effectiveness and NFs in RL. Building on prior work [45, 19], we develop a simple NF architecture and evaluate it across 82 tasks spanning five different RL settings, including offline imitation learning, offline RL, goal-conditioned RL, and unsupervised RL:
  • On offline imitation learning, NFs are competitive with baselines using state of the art generative models such as diffusion and transformers, while using significantly less compute (6$\times$ fewer parameters than diffusion policies).
  • On offline conditional imitation learning, NFs outperform BC baselines using models like flow matching, and offline RL algorithms which learn a value function, across 45 tasks.
  • On offline RL problems, NFs are competitive with strong baselines using diffusion / flow matching models on most tasks, outperforming all of them on $15/30$ tasks.
  • On unsupervised GCRL, NFs can rival contrastive learning based algorithms, despite making fewer assumptions.
2 Related Work
Normalizing flows.
This work builds on the rich literature of normalizing flow architectures [18, 72, 45, 46, 39, 65, 89]. NICE [18] introduce the idea of stacking neural function blocks that perform shifting, making density estimation efficient. RealNVP [19] extends NICE to non volume preserving layers that also perform a scaling operation. GLOW [45] further propose applying an invertible generalized permutation and a special normalization layer (ActNorm) to improve expressivity.
Probabilistic models in RL.
Strides in probabilistic modeling research consistently push the boundaries in deep RL. Variational autoencoders gave rise to a family of model-based [34, 35, 91] and representation learning [51, 88] algorithms. Contrastive learning has been used to learn value functions [58, 73, 24, 61] as well as representations [50, 79] for RL. Advances in sequence modeling led to improvements in BC [13, 10, 52, 74], RL [62, 31] and planning [41]. More recently, diffusion models [78, 37, 76] and their variants [77, 56, 5, 6] have become the popular models for large scale BC pre-training [14, 55, 87, 7], RL finetuning [71, 36], offline RL [2, 67, 85] and planning [40, 90, 57].
NFs in RL.
In RL, NFs have been used to learn policies from uncurated data [75] and model exact MaxEnt policies [11, 32, 11]. But despite their flexibility (see Fig. 1 and Table 1), they have not been go-to models in RL; research exploring the application of NFs in RL remains limited. We believe this is because historically, NFs were thought to have limited expressivity [89, 64]. In our experiments, we show that this is not the case (Section 6). We hope that our work brings greater attention towards the potential of NFs in RL, and the development of scalable probabilistic models with similar desirable properties (Fig. 1).
3 Preliminaries
This section describes two major approaches for training probabilistic models and how NFs can be trained using both of them. In the next section, we show that this flexibility allows NFs to be directly plugged into various RL algorithms, improving their expressivity and performance.
Maximum likelihood estimation (MLE).
Let $x\in\mathbb{R}^{d}$ be a random variable with an unknown density $p(x)$. Given a dataset of samples from this density $\mathcal{D}=\{x_{i}\}_{i=1}^{N}\sim p(x)$, the goal of MLE is to estimate the parameters of a probabilistic model $p_{\theta}(x)$ that maximize the likelihood of the dataset:
${\color[rgb]{0.34765625,0.546875,0.90625}\theta_{\text{MLE}}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\big{[}\log p_{\theta}(x)\big{]}}.$ (1)
If $p_{\theta}(x)$ belongs to a universal family of distributions, then the MLE estimate $p_{\theta_{\text{MLE}}}(x)$ converges towards the true density, as dataset size $N$ increases [42].
Variational inference (VI).
In many problems, instead of samples from the true density $p(x)$, one has access to its unnormalized part $\tilde{p}(x)=p(x)\times Z$. The goal of variational inference is to approximate $p(x)$ using a distribution $q$ from some variational family of distributions ${\mathcal{Q}}$:
${\color[rgb]{0.4765625,0.28515625,0.53515625}q^{*}(x)=\operatorname*{arg\,min}_{q\in\mathcal{Q}}D_{\text{KL}}(q\mid\mid p)=\operatorname*{arg\,max}_{q\in\mathcal{Q}}\mathbb{E}_{x\sim q(x)}\big{[}\log\tilde{p}(x)-\log q(x)\big{]}},$ (2)
where Z can be ignored from the last equation. If $\mathcal{Q}$ is universal, the global optima of Eq. 2 will be the true density $q^{*}(x)=p(x)$. To optimize Eq. 2 using gradient descent, one needs to choose a family $\mathcal{Q}$, from which it is easy generate samples (the expectation requires sampling from $q$), evaluate likelihoods (second term), and backpropagate through both these operations. Due to these constraints, one typically has to restrict $\mathcal{Q}$ to be a simpler family [43] (e.g., exponential family), or resort to other inference algorithms [30] that require additional complexity (more hyper-parameters or models).
Normalizing flows (NFs).
NFs learn an invertible mapping $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, from true density $p(x)$ to the a simple prior density $p_{0}(x)$. This function is parameterized such that the inverse $f_{\theta}^{-1}$ as well as the determinant of its jacobian matrix ($|\frac{df_{\theta}(x)}{dx}|$) is easily calculable. The learned density $p_{\theta}(x)$ can then be estimated using the change of variable formula:
$p_{\theta}(x)=p_{0}(f_{\theta}(x))|\frac{df_{\theta}(x)}{dx}|.$ (3)
To sample from an NF, one first samples from the prior, and return the inverse function:
$z\sim p_{0}(x),\;\text{and return}\;x=f_{\theta}^{-1}(z).$ (4)
Because Eq. 3 and Eq. 4 are passes through a single neural network, back-propagating gradients through them is straightforward. Hence, the NF can be trained efficiently using both MLE (Eq. 1) and VI (Eq. 2[81, 72]:
${\color[rgb]{0.34765625,0.546875,0.90625}\theta_{\text{MLE}}}$ ${\color[rgb]{0.34765625,0.546875,0.90625}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{x\sim\mathcal{D}}\big{[}\log p_{0}(f_{\theta}(x))+\log(|\frac{df_{\theta}(x)}{dx}|)\big{]}}.$ (5)
${\color[rgb]{0.4765625,0.28515625,0.53515625}\theta_{\text{VI}}}$ ${\color[rgb]{0.4765625,0.28515625,0.53515625}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{\begin{subarray}{c}z\sim p_{0}(x)\\ x=f^{-1}_{\theta}(z)\end{subarray}}\big{[}\log\tilde{p}(x)-\log p_{0}(z)+\log(|\frac{df^{-1}_{\theta}(x)}{dx}|)\big{]}}.$ (6)
Equation 5 follows from substituting the change of variable formula in Eq. 1. Equation 6 is derived by the same substitution, followed by noting that $z=f_{\theta}(f^{-1}_{\theta}(z))$ and $-\log(|\frac{df^{-1}_{\theta}(x)}{dx}|)=\log(|\frac{df_{\theta}(x)}{dx}|)$. It is standard practice to stack multiple such mappings together $f_{\theta}=f_{\theta}^{1}\circ f_{\theta}^{1}\circ\dots\circ f_{\theta}^{T}$. Each mapping is referred to as a block in the entire NF.
4 Many RL Algorithms do MLE and VI.
In this section, we first introduce multiple RL problems settings. We will then show how popular representative algorithms in each setting are instantiations of MLE and VI. This will $(1)$ underscore the importance of flexible models that can be directly optimized using MLE and VI, and $(2)$ motivate our use of NFs as a generally capable family of models for RL. We will then describe our architecture in the next section (Section 5). In all the following problems, we assume an underlying controlled markov process $\mathcal{M}$, with states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$. The dynamics are $p(s^{\prime}|s,a)$, the initial state distribution is $p_{0}(s_{0})$ and $H$ is the episode horizon.
4.1 Offline Imitation Learning (IL)
Given a dataset of state-action trajectories $\mathcal{D}=\{s_{0}^{i},a_{0}^{i},\dots,s_{H}^{i}\ \}_{i=1}^{N}$ collected by an expert(s), the aim in IL is to match the data generating behavior without any environment interactions.
Behavior Cloning (BC) [4, 69]
is an algorithm that models the conditional distribution of actions given states directly using MLE:
$\operatorname*{arg\,max}_{\theta}{\color[rgb]{0.34765625,0.546875,0.90625}\underbrace{\mathbb{E}_{s_{t},a_{t}\sim\mathcal{D}}\big{[}\log{\pi}_{\theta}(a_{t}\mid s_{t})\big{]}}_{\text{MLE}}}.$ (7)
Goal Conditioned Behavior Cloning (GCBC) [21, 17, 29]
is a simple variant of BC which uses a goal $g\in\mathcal{S}$ as additional conditioning information. It is based on the insight that a trajectory leading to any goal, can be treated as an optimal trajectory to reach that goal. Given a state-action pair, a state sampled from the future $p_{t+}(s_{t+}\mid s_{t},a_{t})$ is treated as the goal. Formally, GCBC maximizes the following objective:
$\operatorname*{arg\,max}_{\theta}{\color[rgb]{0.34765625,0.546875,0.90625}\underbrace{\mathbb{E}_{\begin{subarray}{c}s_{t},a_{t}\sim\mathcal{D}\\ g_{t+}\sim p(s_{t+}\mid s_{t},a_{t})\end{subarray}}\big{[}\log{\pi}_{\theta}(a_{t}\mid s_{t},g_{t+})\big{]}}_{\text{MLE}}}.$ (8)
Following prior work [66, 23], we set the future time $t+$ to be a truncated geometric distribution (from $t+1$ to $H$).
4.2 Offline RL
Offline RL assumes an MDP $\mathcal{M}$ with a reward function $r(s_{t},a_{t})$. Given a dataset of trajectories $\mathcal{D}=\{s_{0}^{i},a_{0}^{i},r^{i}_{0},\dots,s_{H}^{i},r_{H}^{i}\ \}_{i=1}^{N}$, the goal is to find a policy $\pi_{\theta}(a_{t}\mid s_{t})$, which maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta},p}[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})]$ [54, 48].
MaxEnt RL + BC [27, 53, 33].
We write out a minimal objective for an offline actor-critic algorithm [27]:
$\operatorname*{arg\,min}_{Q_{\phi}}\mathbb{E}_{\begin{subarray}{c}s_{t},a_{t},r_{t},s_{t+1}\sim D\\ a_{t+1}\sim\pi_{\theta}(\cdot\mid s_{t+1})\end{subarray}}\left[\left(Q_{\phi}(s_{t},a_{t})-r_{t}-\gamma Q_{\mkern 1.5mu\overline{\mkern-1.5mu\phi\mkern-1.5mu}\mkern 1.5mu}(s_{t+1},a_{t+1})\right)^{2}\right],$ (9)
$\operatorname*{arg\,max}_{\pi_{\theta}}{\color[rgb]{0.4765625,0.28515625,0.53515625}\underbrace{\mathbb{E}_{s_{t}\sim D,a^{\pi}_{t}\sim\pi_{\theta}(\cdot\mid s_{t})}\big{[}Q_{\phi}(s_{t},a^{\pi}_{t})-\lambda\log\pi_{\theta}(a^{\pi}_{t}\mid s_{t})\big{]}}_{\text{VI}}}+{\color[rgb]{0.34765625,0.546875,0.90625}\underbrace{\alpha\mathbb{E}_{s_{t},a_{t}\sim D}\big{[}\log\pi_{\theta}(a_{t}\mid s_{t})\big{]}}_{\text{MLE}}},$ (10)
where $Q_{\phi}(s,a)$ is the action-value function and $Q_{\mkern 1.5mu\overline{\mkern-1.5mu\phi\mkern-1.5mu}\mkern 1.5mu}(s,a)$ denotes the target network [60], $\lambda$ and $\alpha$ control the strength of the entropy and behavior regularization respectively. For $\lambda=0$, this algorithm exactly matches TD3+BC [27]. In all our experiments, we use $\lambda=0$ for minimality [^2] This is a special case of variational inference. In bayesian inference this corresponds to approximating a maximum a-posteriori estimate.. The policy has to sample actions that have high values while staying close to the (potentially multi-modal) behavior policy. NFs ability to be trained using both MLE and VI allows the use of expressive policies, without any changes to the simplest offline RL algorithms.
4.3 Goal conditioned RL (GCRL)
GCRL [44] is a multi-task RL problem where tasks correspond to reaching goals states $g\in\mathcal{S}$ sampled from a test goal distribution $p_{\text{test}}(g)$. We will define the discounted state occupancy distribution as $p^{\pi}_{t+}(s_{t+}\mid s,a)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s_{t+}\mid s_{0},a_{0}=s,a)$. [^3] We will work with in infinite horizon case, with $H=\infty$.
Q-function estimation.
Prior work has shown that the Q-function for GCRL is equal to the discounted state occupancy distribution $Q^{\pi}(s,a,g)=p^{\pi}_{t+}(s_{t+}\mid s,a)$ [24]. Hence estimating the Q-function can be framed as a maximum likelihood problem:
$\operatorname*{arg\,max}_{\theta}{\color[rgb]{0.34765625,0.546875,0.90625}\underbrace{\mathbb{E}_{\begin{subarray}{c}s_{t},a_{t}\sim\mathcal{D}\\ g_{t+}\sim p_{t+}(s_{t+}\mid s_{t},a_{t})\end{subarray}}\big{[}\log p_{\theta}(g_{t+}\mid s_{t},a_{t})\big{]}}_{\text{MLE}}}.$ (11)
Unsupervised goal sampling (UGS) [25, 68]
algorithms do not assume that the test time goal distribution $p_{\text{test}}(g)$ is known apriori. The main aim of UGS is to maximize exploration by commanding goals that are neither too easy nor too difficult for the current agent. This inclines the agent to explore goals on the edge of its goal coverage. But to recognize which goals are on the edge, one needs an estimate of the goal coverage density $p_{t+}(g)$. Many methods [70, 68, 25] primarily fit a density model and then choose goals that have low density. A natural way of estimating densities is MLE, though some prior works use secondary objectives [63, 80, 25, 70] or non-parametric models [68]:
$\operatorname*{arg\,max}_{\theta}{\color[rgb]{0.34765625,0.546875,0.90625}\underbrace{\mathbb{E}_{\begin{subarray}{c}s_{t},a_{t}\sim\mathcal{D}\\ g_{t+}\sim p_{t+}(s_{t+}\mid s_{t},a_{t})\end{subarray}}\big{[}\log p_{\theta}(g_{t+})\big{]}}_{\text{MLE}}}.$ (12)
Using an estimate of this density, the canonical UGS algorithm is to sample a large batch of goals from the replay buffer and command the agent to reach the goal with the minimum density under $p_{\theta}(g_{t+})$ [68].
5 A simple and efficient NF architecture for general RL
This section will introduce the NF architecture we use and Section 6 will show how the this architecture fits seamlessly into many RL algorithms (those from Section 4). Many design choices affect model expressivity, the speed of computing the forward ($f_{\theta}(x)$) and inverse ($f_{\theta}^{-1}(x)$), and the efficiency of computing the determinant of the Jacobian [64]. Our desiderata is to build a simple and fast architecture that is expressive enough to compete and even surpass the most popular approaches across a range of RL problems. To achieve this, we build each NF block $f^{t}_{\theta}$ by leveraging two components from prior work: (1) a fully connected “coupling” network [19], and (2) a linear flow [45]. In all of our experiments, the model learns a conditional distribution $p_{\theta}(x\mid y)$, so each NF block $f^{t}_{\theta}:\mathbb{R}^{d}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ takes in $y\in\mathbb{R}^{k}$ as an additional input.
Coupling network.
We use an expressive and simple transformation proposed in Dinh et al., [18, 19]. Let $x_{t}$ be the input for the $t^{\text{th}}$ block and $y$ be the conditioning information. For the $t^{\text{th}}$ block, the coupling network $f_{\theta}^{t}=\{a_{\theta}^{t},s_{\theta}^{t}:\mathbb{R}^{\lfloor d/2\rfloor}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{\lceil d/2\rceil}\}$ does the following computation:
$(1)\;x_{1}^{t},x_{2}^{t}=\text{split}(x^{t}),\;(2)\;\tilde{x}_{2}^{t}=(x_{2}^{t}+a_{\theta}^{t}(x_{1}^{t},y))\times\exp(-s_{\theta}^{t}(x_{1}^{t},y)),\;(3)\;\tilde{x}^{t}=\text{concat}(x_{1}^{t},\tilde{x}_{2}^{t}).$
Note that the log-det Jacobian of $f^{t}_{\theta}$ is $-\text{sum}(s_{\theta}^{t}(x_{1}^{t},y))$. Further, given $\tilde{x}^{t}$ and $y$, the inverse ${f^{t}_{\theta}}^{-1}$ can be calculated as easily as the forward map (see Appendix B for details).
Linear flow.
We adapt the generalized permutation proposed by Kingma and Dhariwal, [45] to one dimensional inputs, to linearly transform the output of the coupling layer $\tilde{x}^{t}$. The intuition behind adding this linear flow is that each dimension of $\tilde{x}^{t}$ can be transformed by other dimensions in an end to end manner:
$x^{t+1}=(\tilde{x}^{t})^{T}W,\;\text{where}\;W=PL_{\theta}U_{\theta},\;\text{and}\;\log|W|=\text{sum}\log(|\text{diag}(U)|).$
Here, $P$ is a fixed permutation matrix and $L_{\theta}$ is lower triangular with ones on the diagonal and $U_{\theta}$ is upper triangular. Hence, inverting them incurs similar costs as matrix multiplication [64] and calculating the log-det Jacobian requires only the diagonal elements of $U$.
Normalization.
While previous approaches used special variants of normalization layers ([19], [45]), we find that just using LayerNorm [3] in the coupling network’s weights was sufficient of training deep NFs.
Overall, our architecture can be thought of as RealNVP [19] with generalized permutation, or a linear version of GLOW [45] without ActNorm. We have released a simple code for this architecture both in Jax and PyTorch: code link.
6 Experiments
Figure 2: Visualization of tasks
The main aim of our experiments is to use NFs and question some of the standard assumptions across several RL settings. $(1)$ In IL, methods like diffusion policies [14] are popular because they provide expressivity but at a high computational cost. Our experiments with NFs show that this extra compute might not be necessary (Section 6.1). $(2)$ In conditional IL, prior work argues that learning a value function is necessary for strong performance [49, 12, 28, 1]. Our experiments suggest that a sufficiently expressive policy, such as an NF, can often suffice (Section 6.2). $(3)$ In offline RL, auxiliary techniques are typically used to support expressive policies. We show that NF policies can be directly integrated with existing offline RL algorithms to achieve high performance (Section 6.3). $(4)$ In UGS, we show that a single NF can estimate both the value function and goal density, enabling better exploration compared to baseline that make stronger assumptions.
In all experiments, we use the architecture described in Section 5 to implement all algorithms described in Section 4. Implementation details for all algorithms are provided in Appendix A. Across all problems (IL, offline RL, online GCRL), we present results on a total of 82 tasks (see Fig. 2 for visualization of tasks) with 5 seeds each, which show that NFs are generally capable models that can be effectively applied to a diverse set of RL problems. Unless stated otherwise, all error bars represent $\pm$ standard deviation across seeds.
6.1 Imitation Learning (IL)
Figure 3: Imitation Learning. ($a-d$) Results on multi-modal behavior datasets which require fine grained control and expressivity. Simplicity is measured by the inverse of the number of hyper-parameters (see Appendix D for details) and performance is measured using normalized rewards. NF-BC is the simplest method, using  2 to 3 times less hyper-parameter than VQ-BeT and DiffPolicy, and performs competitively on all tasks. It outperforms DiffPolicy [14] and VQ-BeT [52] on ${2}/{4}$ tasks, ($e$) while using only a fraction of parameters compared to DiffPolicy-C. These results challenge the notion that NFs are inexpressive. We only report standard errors for our methods as baseline results were taken from prior work which did not include error bars [52].
Figure 4: BC baselines architectures
Algorithm Model types
BeT Transformer.
VQ-BeT Transformer + VQ-VAEs.
DiffPolicy-C Convolutional network.
DiffPolicy-T Transformer.
NF-BC Fully connected NFs.
In this section, we use our NF architecture with BC (Section 4.1, Eq. 7) for imitation learning (NF-BC). Please see implementation details of training and sampling with IL algorithm in Section A.1.
Tasks. PushT requires fine-grain control to maneuver a T-block in a fixed location. Performance is measured by the final normalized area of intersection between the block and the goal. Multimodal Ant, UR3 BlockPush, and Kitchen contain multiple behaviors (4, 2, and 4 respectively). Performance is measured by the normalized number of distinct behaviors the agent is able to imitate successfully over 50 independent trials. We also report the simplicity of algorithms, as it is often a decisive factor in determining whether the algorithm can be practically employed. Simplicity is measured by the negating the number of hyperparameters used. We list and briefly describe all the hyperparameters of all the methods used to make Fig. 3 in  Appendix D. Baselines. We compare our algorithms with 4 baselines. Diffusion policies [14] use high performing diffusion model architectures. We use both the convolutional and transformer (DiffPolicy-C and DiffPolicy-T) variant of the diffusion policy. VQ-BeT [52] uses VQ-VAEs [83] for discretizing actions and learns a transformer model over chunked representations. BeT [74] is similar to VQ-BeT but uses K-means for action chunking.
Results. In Fig. 3 we see that NF-BC outperforms all baselines on PushT and Kitchen and performs competitively in other two tasks. Of all the baselines, NF-BC is the simplest by more than a factor of  2. For example, DiffPolicy-T uses 3 times more hyperparameters than NF-BC. These results suggest that NF policies do not lack in expressivity, compared to both diffusion and auto-regressive policies despite introducing fewer additional hyper-parameters. In fact, NF-BC can achieve these results using just a fully connected network and fraction of parameters compared to diffusion policies (see rightmost plot in Fig. 3). Unlike BeT [74] and VQ-BeT [52], NF-BC does not use autoregressive models or VQ-VAEs (see Fig. 4 for a comparison of architectures). Reducing the number of hyperparameters is especially important for practitioners because it makes implementation easier and decreases the need for hyperparameter tuning.
6.2 Conditional Imitation Learning
In this section, we use our NF architecture with GCBC (Section 4.1, Eq. 8) for conditional imitation learning (NF-GCBC).
Tasks. We use a total of 45 tasks from OGBench [66] meant to test diverse capabilities like learning from suboptimal trajectories with high dimensional actions, trajectory stitching and long-horizon reasoning. Performance is measured using the average success rate of reaching a goal in 50 independent trials. Baselines. We compare with 6 different baselines. GCBC is the GCBC 4.1 algorithm with a gaussian policy, and FM-GCBC is the GCBC 4.1 algorithm with a flow-matching based policy [56]. The velocity field for FM-GCBC is a fully connected network trained using flow matching. FM-GCBC is the closest baseline to NF-GCBC as it uses the same underlying algorithm, uses a fully connected network like NF-GCBC, and uses an expressive family of policy models. GCIQL [47] is an offline RL algorithm that learns the optimal value function corresponding to the best actions in the dataset. QRL [86] uses a quasimetric architecture to learn the optimal value function and CRL [24] uses contrastive learning to estimate behavior policy’s value function.
Figure 5: Conditional Imitation Learning. (Left)  NF-GCBC exceeds its gaussian counterpart GCBC on ${40}/{45}$ tasks, demonstrates that using expressive models like NFs can significantly improve performance. (Right)  Despite being a BC–based algorithm, NF-GCBC outperforms all offline RL baselines across 45 tasks. NF-GCBC performs 77% better than FM-GCBC which is the closest baseline that uses a flow matching policy with GCBC.
Results. In Fig. 5 (left), we see that NF-GCBC exceeds the performance (often significantly) of GCBC on ${40}/{45}$ tasks. This underscores the importance of using an expressive policy with the GCBC algorithm. In Fig. 5 (right), NF-GCBC, despite being a BC based algorithm, outperforms all baselines, including dedicated offline RL algorithms. We have added separate task wise plots in Appendix C. NF-GCBC outperforms FM-GCBC, which is the closest baseline, by 77%. While this does not necessarily imply that NFs are more expressive than flow-matching models, we hypothesize that all else being equal (e.g., fully connected architectures and identical underlying algorithms), NFs are better able to imitate the data using simple architectures as they directly optimize for log likelihoods. Flow matching models on the other hand estimate probabilities indirectly (by learning an SDE / ODE), and might require more complex architectures (transformers or convolutional networks) to work well.
6.3 Offline RL
Figure 6: Offline RL. NF-RLBC outperforms all baselines on ${15}/{30}$ tasks, and achieves top 2 results on ${10}/{30}$ tasks. Notably, NF-RLBC achieves much higher results than all baselines on puzzle-3x3-play and scene-play, which require the most long horizon and sequential reasoning. For example, on puzzle-3x3-play NF-RLBC performs 230% better than the next best baseline.
While NF-GCBC performs better than offline RL algorithms on average, there are certain tasks like puzzle-3x3-play where only TD based methods succeed (Appendix C). This task, and many others as we will see, require learning a value function and accurately extracting behaviors from it. Fortunately, NFs can be trained using both VI and MLE (3). This allows us to train MaxEnt RL+BC (Section 4.2, Eq. 9) using our NF architecture as a policy (NF-RLBC).
Tasks. We use a total of 30 tasks from prior work [67] testing offline RL with expressive policies. Each task has a goal which requires completing $K$ subtasks. At each step, the agent gets a reward equal to the negative of the number of subtasks left to be completed. For our experiments, we choose tasks that cover a diverse set of challenges and robots. Tasks like humanoidmaze-medium-navigate and antmaze-medium-navigate contain diverse suboptimal trajectories and high dimensional actions. Antsoccer-arena-navigate requires controlling a quadruped agent to dribble a ball to a goal location. Scene-play requires manipulating multiple objects and puzzle-3x3-play requires combinatorial generalization and long-horizon reasoning. We hypothesize that the Q function for the some of these tasks (esp manipulation tasks like puzzle and scene which require long horizon reasoning) can have narrow modes of good actions and a large number of bad actions. Hence using direct gradient based optimization can be crucial to search for good actions. Baselines. We choose 5 representative baselines. BC performs BC with a gaussian policy, ReBRAC [82] is an offline RL algorithm with a gaussian policy that achieves impressive results on prior benchmarks [26]. IDQL [36] uses importance resampling to improve over a diffusion policy. IFQL is the flow counterpart of IDQL [67]. FQL [67] is a strong offline RL algorithm that distills a flow-matching policy into a one-step policy, while jointly maximizing the Q-function. Notably, all offline RL algorithms (IDQL, IFQL, FQL) have to incorporate additional components—such as distillation or resampling—to leverage policy expressiveness.
Results. Figure 6, shows that on average NF-RLBC outperforms all baselines on ${15}/{30}$ tasks, and achieves top 2 results on ${10}/{30}$ tasks. Notably, NF-RLBC achieves much higher results than all baselines on puzzle-3x3-play and scene-play, which require most long horizon and sequential reasoning (on puzzle-3x3-play NF-RLBC performs 230% better than the next best baseline). NF-RLBC is able to search for good actions because it directly uses gradient based optimization to backpropagate action gradients through the policy. All other offline RL algorithms with expressive policy classes (IDQL, IFQL, FQL) have to use extra intermediate steps.
Figure 7: GCRL. (first 3 from left) NF-UGS consistently outperforms other unsupervised baselines. (rightmost) NF-UGS achieves higher asymptotic success rates than the CRL-oracle averaged across all tasks. Note that CRL-oracle makes extra assumptions about the availability of test-time goals.
6.4 Goal Conditioned RL
In this section, we use our NF architecture to estimate both the Q-function (Eq. 11) and the coverage density (Eq. 12). Both quantities model the distribution over the same random variable (future goals), with the Q-function conditioned on a state-action pair. We use a single neural network to estimate both by employing a masking scheme to indicate the presence or absence of conditioning [38]. We then train a goal-conditioned policy to maximize the Q-function in a standard actor-critic setup, and sample training goals using the canonical UGS algorithm (Section 4.3), resulting in our NF-UGS method.
Tasks. We use three problems from the UGS literature [68, 9] ant-u-maze, ant-big-maze and ant-hard-maze. These environments are difficult because the goal test distribution is unknown, and success requires thorough exploration of the entire maze. Baselines. We compare against three baselines. CRL-oracle [24, 9] uses contrastive learning to estimate the Q-function and serves as an oracle baseline as it uses test goals $p_{\text{test}}(g)$ to collect data during training. CRL-uniform is a variant that samples goals uniformly from the replay buffer, while CRL-minmax selects goals with the lowest Q-value from the initial state of the MDP – intuitively targeting the goals the agent believes are hardest to reach. The aim of these comparisons isn’t to propose a state of the art method, but rather to check whether NFs can efficiently estimate the coverage densities of the non-stationary RL data for unsupervised exploration. Results. These experiments aim to evaluate whether NFs can efficiently estimate both the Q-function and the coverage density. As shown in Fig. 7, NF-UGS outperforms all baselines, and notably the oracle algorithm as well (aggregated score is the average of the final success rates across all three tasks). Unlike the minmax strategy, which plateaus over time, NF-UGS continues to improve by sampling goals to maximize the entropy of the coverage distribution [68].
7 Conclusion
In this paper, we view RL problems from a probabilistic perspective, asking: which probabilistic models provide the capabilities necessary for efficient RL? We argue that NFs offer a compelling answer. NFs outperform strong baselines across diverse RL settings, often with fewer parameters and lower compute, leading to simpler algorithms. By making density estimation simpler and effective, we hope our work opens new avenues for areas such as distributional RL, bayesian RL, and RL safety.
A primary limitation of NFs is that they have restrictive architectures, in that they have to be invertible. Nevertheless, there is a rich line of work characterizing which NF architectures are universal [20, 16, 64]. Another limitation of our work is that we do not explore new NF architectures; the design we use is mainly adapted from prior work (Section 5).
Acknowledgments and Disclosure of Funding
The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Office of Information Technology’s Research Computing. The authors thank Emmanuel Bengio, Liv d’Aliberti and Eva Yi Xie for their detailed reviews on our initial drafts. The authors thank Seohong Park for providing code for some baselines and Catherine Ji for helping find a bug in our code. The authors also thank the members of Princeton RL lab for helpful discussions.
Appendix A Implementation details.
In this section we provide all the implementation details as well as hyper-parameters used for all the algorithms in our experiments – NF-BC, NF-GCBC, NF-RLBC, and NF-UGS. Unless specified otherwise, we chose common hyper-parameters used in prior work without tuning.
A.1 Imitation Learning: NF-BC
For all imitation learning, we use a policy with 12 NF blocks. The fully connected networks of the coupling network for each NF block consist of two hidden layers with 512 activations each. We use a fully connected network for the state-goal encoder with 4 hidden layers having 512 activations each.
It is common in the NF literature to add some noise to the training inputs before training an NF using MLE [72, 89]. For NF-BC, we add gaussian noise to the actions before training. Hence, NF-BC models the noisy behavior policy instead $\pi^{\text{a+noise}}(a\mid s)$. To reduce noise from the obtained samples,  Zhai et al., [89] propose a denoising trick:
$z\sim p_{0},\;y:=f^{-1}(z)\;,x:=y+\sigma^{2}\nabla_{y}\log p_{\theta}(x)$
We use this trick in all our NF-BC experiments. Table 2 contains the values of hyperparameters used by NF-BC. All experiments were done on RTX 3090s, A5000s or A6000s and did not require more than 20 hours. Experiment speed is mostly bottlenecked by intermediate evaluation of BC policies because the environments are slow and the number of rollouts per evaluation is 50. For example, the experiments on the multimodal-ant environment took 20 hours, but on pusht they were finished in less than 90 minutes.
Table 2: NF-BC: Hyper-parameters and their values.
Hyper-parameter Description
channels 512
noise std 0.1
blocks 12
encoder num layers 4
rep dims 512
A.2 Conditional Imitation Learning: NF-GCBC
For all conditional imitation learning experiments, we use a policy with 6 NF blocks. The fully connected networks of the coupling network for each NF block consist of two hidden layers with 256 activations each. We use a fully connected network for the state-goal encoder with 4 hidden layers having 512 activations each. The output representation dimensions are 64. We also employ the denoising trick proposed by [89] and mentioned in Section A.1. Table 3 contains the values of hyperparamters used by NF-GCBC. All experiments were done on RTX 3090s, A5000s or A6000s and did not require more than 3 hours each.
Table 3: NF-GCBC: Hyper-parameters and their values.
Hyper-parameter Description
channels 512
noise std 0.1
blocks 6
encoder num layers 4
rep dims 512
Future goal sampling discount Table 4
Table 4: NF-GCBC: Environment values for future goal sampling discount values. For all GCBC algorithms, including the GCBC baselines, we tuned this value in the set {0.97, 0.99}
Environment Name Future goals sampling discount value $\gamma$
puzzle $3\times 3$ play 0.97
scene play 0.97
cube single play 0.97
cube double play 0.99
humanoidmaze medium navigation 0.97
antmaze large navigation 0.99
antsoccer arena navigation 0.97
antmaze medium stitch 0.97
antsoccer arena stitch 0.97
A.3 Offline RL: NF-RLBC
For all offline RL experiments, we use a policy with 6 NF blocks. The fully connected networks of the coupling network for each NF block consist of two hidden layers with 256 activations each. We use a fully connected network for the state-goal encoder with 4 hidden layers having 512 activations each. The output representation dimensions are 64. Table 5 contains the values of main hyperparameters used by NF-RLBC. All experiments were done on RTX 3090s, A5000s or A6000s and did not require more than 3 hours each.
Table 5: NF-RLBC: Hyper-parameters and their values.
Hyper-parameter Description
channels 512
noise std 0.1
actor entropy coefficient ($\lambda$ in Eq. 9) 0.0
actor BC loss coefficient ($\alpha$ in Eq. 9) Table 6
blocks 6
encoder num layers 4
rep dims 512
Table 6: NF-RLBC: Environment wise actor BC loss coefficient. We use the same hyperparameter tuning procedure as all the baselines [67]. See App E.2. of Park et al., 2025b [67]
Environment Name Actor BC loss coefficient $\alpha$
puzzle $3\times 3$ play 10.0
scene play 10.0
antmaze large navigation 1.0
cube single play 10.0
humanoidmaze medium navigation 1.0
antsoccer arena navigation 1.0
A.4 Unsupervised RL: NF-UGS
For all unsupervised RL experiments, we use a goal conditioned value function with 6 NF blocks. The fully connected networks of the coupling network for each NF block consist of two hidden layers with 256 activations each. We use a fully connected network for the state-action encoder with 4 hidden layers having 1024 activations each. We train both the conditional $p_{\theta}(g\mid s,a)$ and the unconditional $p_{\theta}(g)$ using the same network. We apply a mask on the conditional variable inputs with probability $0.1$. The discount factor we use is $0.99$. We add a gaussian noise with standard deviation $0.05$. Because we do not sample from the model in NF-UGS, we do not need to employ the denoising trick. Instead, 1024 random goals from the buffer are sampled and the one with the lowest marginal density $p_{\theta}(g)$ is selected for data collection. All experiments were done on RTX 3090s, A5000s or A6000s and did not require more than 20 hours each.
Appendix B More details about NF block.
Figure 8: NF block.
It requires the same cost as the coupling of the forward block to calculate the coupling of the inverse block ${f^{t}_{\theta}}^{-1}$:
$\tilde{x}_{1}^{t},\tilde{x}_{2}^{t}=\text{split}(\tilde{x}^{t}),$
$x_{1}^{t}=\tilde{x}_{1}^{t},$
$x_{2}^{t}=\tilde{x}_{2}^{t}\times\exp(s_{\theta}^{t}(x_{1}^{t},y))-a_{\theta}^{t}(x_{1}^{t},y)),$
$x^{t}=\text{concat}(x_{1}^{t},x_{2}^{t}).$
See Fig. 8 for a visualization of each NF block.
Appendix C Additional results.
Table 7 and Table 8 contain additional experimental results.
Table 7: Task wise performance for Fig. 5.
Environment Score
puzzle-3x3-play-v0 $0.045\scriptstyle{\pm 0.005}$
cube-single-play-v0 $0.795\scriptstyle{\pm 0.109}$
scene-play-v0 $0.420\scriptstyle{\pm 0.008}$
cube-double-play-v0 $0.351\scriptstyle{\pm 0.087}$
antmaze-medium-stitch-v0 $0.566\scriptstyle{\pm 0.068}$
humanoidmaze-medium-navigate-v0 $0.314\scriptstyle{\pm 0.042}$
antmaze-large-navigate-v0 $0.320\scriptstyle{\pm 0.055}$
antsoccer-arena-navigate-v0 $0.621\scriptstyle{\pm 0.007}$
antsoccer-arena-stitch-v0 $0.643\scriptstyle{\pm 0.016}$
[…]
Environment Score
antmaze-large-navigate-singletask-task1-v0 $0.764\scriptstyle{\pm 0.073}$
antmaze-large-navigate-singletask-task2-v0 $0.724\scriptstyle{\pm 0.088}$
antmaze-large-navigate-singletask-task3-v0 $0.924\scriptstyle{\pm 0.023}$
antmaze-large-navigate-singletask-task4-v0 $0.616\scriptstyle{\pm 0.314}$
antmaze-large-navigate-singletask-task5-v0 $0.608\scriptstyle{\pm 0.399}$
antsoccer-arena-navigate-singletask-task1-v0 $0.600\scriptstyle{\pm 0.122}$
antsoccer-arena-navigate-singletask-task2-v0 $0.556\scriptstyle{\pm 0.096}$
antsoccer-arena-navigate-singletask-task3-v0 $0.420\scriptstyle{\pm 0.077}$
antsoccer-arena-navigate-singletask-task4-v0 $0.480\scriptstyle{\pm 0.061}$
antsoccer-arena-navigate-singletask-task5-v0 $0.336\scriptstyle{\pm 0.062}$
cube-single-play-singletask-task1-v0 $0.440\scriptstyle{\pm 0.057}$
cube-single-play-singletask-task2-v0 $0.640\scriptstyle{\pm 0.087}$
cube-single-play-singletask-task3-v0 $0.768\scriptstyle{\pm 0.060}$
cube-single-play-singletask-task4-v0 $0.600\scriptstyle{\pm 0.082}$
cube-single-play-singletask-task5-v0 $0.616\scriptstyle{\pm 0.067}$
humanoidmaze-medium-navigate-singletask-task1-v0 $0.036\scriptstyle{\pm 0.045}$
humanoidmaze-medium-navigate-singletask-task2-v0 $0.188\scriptstyle{\pm 0.097}$
humanoidmaze-medium-navigate-singletask-task3-v0 $0.148\scriptstyle{\pm 0.072}$
humanoidmaze-medium-navigate-singletask-task4-v0 $0.016\scriptstyle{\pm 0.015}$
humanoidmaze-medium-navigate-singletask-task5-v0 $0.208\scriptstyle{\pm 0.053}$
puzzle-3x3-play-singletask-task1-v0 $0.996\scriptstyle{\pm 0.008}$
puzzle-3x3-play-singletask-task2-v0 $0.996\scriptstyle{\pm 0.008}$
puzzle-3x3-play-singletask-task3-v0 $0.976\scriptstyle{\pm 0.029}$
puzzle-3x3-play-singletask-task4-v0 $0.996\scriptstyle{\pm 0.008}$
puzzle-3x3-play-singletask-task5-v0 $0.996\scriptstyle{\pm 0.008}$
scene-play-singletask-task1-v0 $1.000\scriptstyle{\pm 0.000}$
scene-play-singletask-task2-v0 $0.880\scriptstyle{\pm 0.075}$
scene-play-singletask-task3-v0 $0.988\scriptstyle{\pm 0.016}$
scene-play-singletask-task4-v0 $0.856\scriptstyle{\pm 0.056}$
scene-play-singletask-task5-v0 $0.000\scriptstyle{\pm 0.000}$
Appendix D Hyperparameters used by behavior cloning algorithms based on different probabilistic models.
Table 9: Hyper-parameters used by NF-BC.
Hyper-parameter Description
channels width of each layer of the coupling MLP network.
noise std added noise to labels for training the NF using MLE.
blocks number of NF blocks.
encoder num layers number of layers in the observation encoder MLP network.
rep dims output dimension of observation encoder MLP network.
Table 10: Hyper-parameters used by BET.
Hyper-parameter Description
embedding dims dimension length for attention vectors.
num layers number of attention blocks.
num heads number of heads in attention.
act scale action scaling parameter.
num bins number of centroids for k-means.
offset prediction whether to predict offsets for actions or not.
offset loss scale action offset prediction loss coefficient.
gamma focal loss parameter.
Table 11: Hyper-parameters used by DiffPolicy-C.
Hyper-parameter Description
time-step embedding dims embedding dimension of diffusion timestep.
down dims downsampling dimensions for Unet architecture.
kernel size kernel size for CNNs in Unet.
n groups number of groups for group normalization.
scheduler train timesteps Length of the forward diffusion chain while training.
num inference steps Length of the denoising steps while sampling.
beta schedule type type of scheduler for noising.
beta start first-step noise for diffusion scheduler.
beta end last-step noise for diffusion scheduler.
ema inverse gamma parameter for ema rate schedule.
ema power parameter for ema rate schedule.
ema min minimum value of ema decay rate.
ema max maximum value of ema decay rate.
Table 12: Hyper-parameters used by DiffPolicy-T.
Hyper-parameter Description
time-step embedding dims embedding dimension of diffusion timestep.
embedding dims dimension length for attention vectors.
num layers number of attention blocks.
num heads number of heads in attention.
scheduler train timesteps Length of the forward diffusion chain while training.
num inference steps Length of the denoising steps while sampling.
beta schedule type type of scheduler for noising.
beta start first-step noise for diffusion scheduler.
beta end last-step noise for diffusion scheduler.
ema inverse gamma parameter for ema rate schedule.
ema power parameter for ema rate schedule.
ema min minimum value of ema decay rate.
ema max maximum value of ema decay rate.
Table 13: Hyper-parameters used by VQ-BET.
Hyper-parameter Description
embedding dims dimension length for attention vectors.
num layers number of attention blocks.
num heads number of heads in attention.
action seq len sequence length used to train the behavior transformer.
vqvae num embed cookbook size of vqvae.
vqvae embedding dims embedding dimensions of vqvae codes.
vqvae groups number of groups in residual vqvae.
encoder loss multiplier encoder loss coefficient.
offset loss multiplier action offset prediction loss coefficient.
secondary code multiplier secondary code prediction loss coefficient.
gamma focal loss parameter.

### v3

@@
 This paper demonstrates that neither of these notions hold NFs back in practice: NFs can match the expressivity of diffusion and autoregressive models while being significantly simpler.
-Moreover, unlike other models, NFs do not require additional machinery to be integrated with existing RL frameworks, enabling simpler algorithms that are competitive with if not better than their counterparts.
+Moreover, unlike other models, NFs do not require additional machinery to be integrated with existing RL frameworks, enabling simpler algorithms that are competitive with, if not better than their counterparts.
 Our primary contribution is to shed light on the effectiveness and NFs in RL.
-Building on prior work [45, 19], we develop a simple NF architecture and evaluate it across 82 tasks spanning five different RL settings, including offline imitation learning, offline RL, goal-conditioned RL, and unsupervised RL:
+Building on prior work [45, 19, 89], we develop a simple NF architecture and evaluate it across 82 tasks spanning five different RL settings, including offline imitation learning, offline RL, goal-conditioned RL, and unsupervised RL:
   • On offline imitation learning, NFs are competitive with baselines using state of the art generative models such as diffusion and transformers, while using significantly less compute (6$\times$ fewer parameters than diffusion policies).
@@
 GLOW [45] further propose applying an invertible generalized permutation and a special normalization layer (ActNorm) to improve expressivity.
+Recently, Zhai et al., [89] propose a scalable NF architecture and a sampling technique to achieve SOTA results on vision benchmarks.
 Probabilistic models in RL.
@@
 Hence, inverting them incurs similar costs as matrix multiplication [64] and calculating the log-det Jacobian requires only the diagonal elements of $U$.
-Normalization.
-While previous approaches used special variants of normalization layers ([19], [45]), we find that just using LayerNorm [3] in the coupling network’s weights was sufficient of training deep NFs.
+Other details.
+While previous approaches used special variants of normalization layers ([19, 45]), we find that just using LayerNorm [3] in the coupling network’s weights was sufficient of training deep NFs.
+For NFs trained purely using MLE (Section 6.1, Section 6.2), we found the noising and denoising technique proposed by [89] (see Section A.1 for details) to be crucial for good performance.
 Overall, our architecture can be thought of as RealNVP [19] with generalized permutation, or a linear version of GLOW [45] without ActNorm.
@@
 $(4)$ In UGS, we show that a single NF can estimate both the value function and goal density, enabling better exploration compared to baseline that make stronger assumptions.
-In all experiments, we use the architecture described in Section 5 to implement all algorithms described in Section 4.
-Implementation details for all algorithms are provided in Appendix A.
+In all experiments, we use the architecture described in Section 5 to implement all algorithms from Section 4, with further implementation details in Appendix A.
 Across all problems (IL, offline RL, online GCRL), we present results on a total of 82 tasks (see Fig.
-2 for visualization of tasks) with 5 seeds each, which show that NFs are generally capable models that can be effectively applied to a diverse set of RL problems.
+2) with 5 seeds each, showing that NFs are generally capable models applicable to diverse RL problems.
 Unless stated otherwise, all error bars represent $\pm$ standard deviation across seeds.
@@
 The authors thank Seohong Park for providing code for some baselines and Catherine Ji for helping find a bug in our code.
-The authors also thank the members of Princeton RL lab for helpful discussions.
+The authors also thank Shuangfei Zhai and the members of Princeton RL lab for helpful discussions.
 Appendix A Implementation details.